# Gauge-dependence of Scalar Induced Gravitational Waves

### v1

@@ Abstract
In this review we look into the gauge-dependence of scalar-induced gravitational waves (SIGWs) that are second-order tensors produced by first-order scalar-modes. The method includes deriving the background, first- and second-order Einstein field equations without imposing a gauge. We address the gauge-invariant approach and study the source-term of SIGWs in three different gauges, synchronous, Poisson and uniform curvature gauge. We find that numerically computed kernels in all three gauges behave closely with minimal discrepancy. As expected, when going in sub-horizon modes, $x\gg 1$, the discrepancy decreases and the behavior matches, pointing to a gauge-invariant observable.
@@ 1 Introduction
Due to the weak interaction with matter, we anticipate to detect primordial gravitational waves (GWs) with current- and future-generation detectors [1]. Sources of GWs from the early universe have been extensively studied in literature [2, 3, 4] and of the various sources one contribution arises from “scalar-induced” GWs (SIGWs) [5, 6, 7, 8, 9, 10]. These are observed as a second-order source of GWs produced by the coupling of large scalar fluctuations at horizon re-entry [11, 12]. The frequency range of SIGWs produced in the radiation dominated era fall within current and future generation ground- and space-based detectors, for example Laser Interferometer Space Antenna (LISA) [13, 14] and Pulsar Timing Array (PTA) [15, 16].
@@
Detailed calculations has been done in previous studies presenting the gauge-dependence of induced GWs by studying different gauges [20, 21, 22]. There have been various ways in approaching this issue. One approach has been by constructing a gauge-invariant formalism for the second-order tensor modes by the addition of terms with the intention to reduce the gauge dependence [23, 24, 25, 26]. However there are some complications with this approach, as pointed out by [27, 28], regarding the understanding between the observable and the gauge-invariant variable. They mention that there is not much difference between choosing particular gauge from the beginning and working with specific gauge-invariant combination.
@@
Instead, the second approach in tackling the gauge issue is by defining a gauge suitable to best describe the observable [27, 29, 28]. In [27], they argue that the suitable gauge is one in which the coordinates are fixed according to the positions of the interferometer mirrors. This is known as the synchronous gauge, or TT gauge. [27] show that the observable GW spectrum of GWs in the TT gauge and Poisson gauge coincide. They discuss that this is a result of the GWs generated during the radiation-dominated era being treated as linear perturbations of the metric as they are freely propagating inside the horizon. Therefore the gauge-dependence does not remain. A similar argument was made by [28], in which they show that the scalar-induced gravitational waves behave similar when calculated in gauges which are “well-behaved” on small scales. The authors point out that the Newtonian gauge is suitable both for the calculation and the physical interpretation. Further mentioning other gauges, among them the synchronous gauge, to also be “well-behaved” in small scales in a universe dominated by perfect fluid with constant equations of state.
@@
In this work, we review the derivation of the source term of scalar-induced gravitational waves in a generic gauge and further on revisit the gauge issue. In Sections 2 and 3, we introduce the general framework for cosmological perturbations of the metric and then derive the Einstein field equations. In Section 3.3, the source term is presented without imposing a gauge. Further on the gauge transformation and gauge-invariant variables are presented in Section 4. We work with the approaches mentioned to show the behaviour of the scalar-induced gravitational waves in different gauges. In Sections 5 and 6, we study the source-term of the SIGWs and the subsequent kernel in three different gauges.
@@ 2 Cosmological perturbation of the metric
Any tensorial quantity, $\mathbf{A}$, can be decomposed into background and inhomogeneous parts,
@@
where $\delta g_{\mu\nu}$ and $\delta^{(2)}g_{\mu\nu}$ is the perturbed metric in first and second order. Using a method known as (3+1) decomposition, 3-dimensional quantities are split into scalar, vector and tensor components corresponding to the transformational behavior on spatial hypersurfaces. The components of the perturbed spatially flat Robertson - Walker metric can be written as
@@
The scalar components are thus $\psi,\phi$, $w$ and $h$, the vectors $w_{i}$ and $h_{i}$ are transverse, i.e. $\partial^{i}w_{i}^{(r)}=\partial^{i}h_{i}^{(r)}=0$, and $h_{ij}^{(r)}$ is a symmetric transverse and trace-free tensor, i.e. $\partial^{i}h_{ij}^{(r)}=h_{i}^{i(r)}=0$. Finally, $D_{ij}=\partial_{i}\partial_{j}-\frac{1}{3}\nabla^{2}\delta_{ij}$ is the trace-free operator.
@@
$g^{00}$ $=-a^{2}(\tau)\left(1-2\phi^{(1)}-\phi^{(2)}+4\left(\phi^{(1)}\right)^{2}-\partial^{i}w^{(1)}\partial_{i}w^{(1)}\right),$
$g^{0i}$ $=a^{2}(\tau)\left[\partial^{i}w^{(1)}+\frac{1}{2}\left(\partial^{i}w^{(2)}+w^{i(2)}\right)+2\left(\psi^{(1)}-\phi^{(1)}\right)\partial^{i}w^{(1)}\right.$
@@
$g^{ij}$ $=a^{2}(\tau)\left[\left(1+2\psi^{(1)}+\psi^{(2)}+4\left(\psi^{(1)}\right)^{2}\right)\delta^{ij}+D^{ij}\left(h^{(1)}+\frac{1}{2}h^{(2)}\right)\right.$
@@
The entire perturbed metric line-element is
$ds^{2}=a^{2}(\tau)$ $\left[-\left(1+2\phi^{(1)}+\phi^{(2)}\right)d\tau^{2}+\left(\partial_{i}w^{(1)}+\frac{1}{2}\partial_{i}w^{(2)}+\frac{1}{2}w_{i}^{(2)}\right)d\tau dx^{i}\right.$
$\left.\left\{\left(1-2\psi^{(1)}-\psi^{(2)}\right)\delta_{ij}+D_{ij}\left(h^{(1)}+\frac{1}{2}h^{(2)}\right)\right.\right.$
$\left.\left.+\frac{1}{2}\left(\partial_{i}h_{j}^{(2)}+\partial_{j}h_{i}^{(2)}+h_{ij}^{(2)}\right)\right\}dx^{i}dx^{j}\right].$ (9)
@@ 3 Einstein field equations
The Einstein field equation is defined as
$G^{\mu}_{\nu}=\kappa^{2}T^{\mu}_{\nu}.$ (10)
@@ 3.1 Background Einsteins equations
where $\bar{P}$ and $\bar{\rho}$ are the background quantities for pressure and energy-density and $\mathcal{H}=\frac{a^{\prime}}{a}$ is the conformal Hubble parameter.
@@ 3.2 First order Einsteins equations
$a^{-2}\left[6\mathcal{H}^{2}\phi^{(1)}+6\mathcal{H}\psi^{(1)^{\prime}}+2\mathcal{H}^{2}\nabla^{2}w^{(1)}-2\nabla^{2}\psi^{(1)}-\frac{1}{2}\partial_{k}\partial^{i}D^{k}_{i}h^{(1)}\right]$ $=-\kappa^{2}\rho^{(1)},$ (12)
$\rho^{(1)}=-\frac{1}{\kappa^{2}a^{2}}\left[6\mathcal{H}\left(\mathcal{H}\phi^{(1)}+\psi^{(1)^{\prime}}\right)+2\mathcal{H}^{2}\nabla^{2}w^{(1)}-2\nabla^{2}\psi^{(1)}-\frac{1}{2}\partial_{k}\partial^{i}D^{k}_{i}h^{(1)}\right].$ (13)
@@
$v_{i}^{(1)}=-\frac{2}{\kappa^{2}a^{2}\left(\bar{\rho}+\bar{P}\right)}\left(\mathcal{H}\partial_{i}\phi^{(1)}+\partial_{i}\psi^{(1)^{\prime}}+\frac{1}{4}\partial_{k}D^{k}_{i}h^{(1)^{\prime}}\right)-\partial_{i}w^{(1)}.$ (15)
@@
$-a^{-2}$ $\left(2\mathcal{H}\phi^{(1)^{\prime}}+4\left[\mathcal{H}^{\prime}+\mathcal{H}^{2}\right]\phi^{(1)}-2\mathcal{H}^{2}\phi^{(1)}+4\mathcal{H}\psi^{(1)^{\prime}}\right.$
$\left.+2\psi^{(1)^{\prime\prime}}+2\nabla^{2}\left(\phi^{(1)}-\psi^{(1)}+2w^{(1)}+w^{(1)^{\prime}}\right)\right)=\kappa^{2}P^{(1)},$ (16)
@@
where we can thus define the anisotropic stress as,
$\pi^{(1)i}_{j}=-\frac{1}{\kappa^{2}}\left(\partial^{i}\partial_{j}-\frac{1}{3}\nabla^{2}\delta^{i}_{j}\right)\left(\phi^{(1)}-\psi^{(1)}+2\mathcal{H}w^{(1)}+w^{(1)^{\prime}}\right).$ (18)
@@ 3.3 Second-order Einstein field equation and source term
For understanding the source-term of scalar-induced gravitational waves, we will be studying the spatial component of second-order Einstein tensor, specifically the transverse, trace-free part. We do this by first applying the projection tensor $P^{li}_{jm}$ to the field equation [30], i.e. $P^{li}_{jm}G^{(2)m}_{l}=\kappa^{2}P^{li}_{jm}T^{(2)m}_{l}$. Accordingly, the second-order vector and tensor modes are removed. Hence the spatial Einstein tensor, Section C.3, and energy-momentum tensor, Section D.3, can be written as
$\left.\left.+2\mathcal{H}w^{(1)}+w^{(1)^{\prime}}\right)\right].$ (20)
Re-arranging and using the relation for the equation of state, $w=\bar{P}/\bar{\rho}$, and the background Section 3.1, we can find the wave-equation for the second-order tensor mode,
@@
where the source term is given in Appendix E. We decided to find this generic form of the source term Appendix E as this can be simplified to further understand the behavior in different gauges by imposing the gauge condition.
@@
where $S_{\lambda}(\tilde{\tau},\mathbf{k})$ includes the Fourier transformation of the source Appendix E and depends on the evolution of the scalar modes. Any scalar mode can be split into a transfer function $T(k\tau)$, which includes the time evolution of the modes after they re-enter the horizon, and the initial condition $\zeta_{\mathbf{k}}$, determined by inflation. For example, the scalar potential $\phi^{(1)}(\mathbf{k},\tau)$ can be split as
$\phi^{(1)}(\mathbf{k},\tau)=\frac{3(1+w)}{5+3w}T_{\phi}(k\tau)\zeta_{\mathbf{k}}\overset{w\rightarrow 1/3}{=}\frac{2}{3}T_{\phi}(k\tau)\zeta_{\mathbf{k}}.$ (24)
@@
where the Green’s function $G_{\mathbf{k}}$ in radiation era is
$G_{\mathbf{k}}(\tau,\tilde{\tau})=\frac{1}{k}\sin(k\tau-k\tilde{\tau})\ .$ (30)
The fraction of GW energy density per logarithmic wavelength is defined by
@@
$\rho_{GW}(\tau)=\frac{1}{16a^{2}\kappa^{2}(\tau)}\overline{\langle(\nabla h_{ij})^{2}\rangle}.$ (32)
The overline above the power spectrum is the oscillation average, derived from the average of the kernel.

### v2

@@ Abstract
In this review we look into the gauge-dependence of scalar-induced gravitational waves (SIGWs) that are second-order tensors produced by first-order scalar modes. The method includes deriving the background, first- and second-order Einstein field equations without imposing a gauge. We address the gauge-invariant approach and study the source-term of SIGWs in three different gauges: synchronous, Poisson, and uniform curvature gauge. We find that numerically computed kernels evaluated in a radiation epoch in all three gauges behave closely with minimal discrepancy. As expected, when going to sub-horizon modes, $k\tau\gg 1$, the discrepancy decreases and the behavior converges, suggesting that SIGWs can be treated as gauge-invariant observables in this regime.
@@ 1 Introduction
Inflationary cosmology leads to the generation of quantum vacuum fluctuations in the metric, giving rise to scalar perturbations—observable today as density fluctuations—and tensor perturbations, which manifest as gravitational waves [1, 2]. These primordial gravitational waves (GWs) are expected to form a stochastic gravitational wave background (SGWB), reflecting their quantum origin. Several early Universe processes act as sources of GWs [3, 4, 5, 6, 7, 8], one of those from scalar-induced gravitational waves (SIGWs) [9, 10, 11, 12, 13, 14]. These are observed as a second-order source of GWs which arise from the nonlinear coupling of large scalar fluctuations at horizon re-entry [15, 16]  [^1] Here and throughout, we refer to ‘second-order’ as perturbations of order $\mathcal{O}(\epsilon^{2})$ arising from products of first-order scalar quantities $\mathcal{O}(\epsilon)$.. The frequency range of SIGWs produced in the radiation dominated era fall within current and future generation ground- and space-based detectors, for example Laser Interferometer Space Antenna (LISA) [17, 18] and Pulsar Timing Array (PTA) [19, 20].
@@
Physically, this raises the question: do SIGWs have any residual gauge dependence at second order? Detailed calculations has been done in previous studies presenting the gauge-dependence of induced GWs by studying different gauges [24, 25, 26]. Some works aim to bypass this issue by constructing a gauge-invariant formalism for the second-order tensor modes by the addition of terms with the intention to reduce the gauge dependence [27, 28, 29, 30]. However, these constructions can blur the connection between theoretical quantities and actual observables [31, 32]. They mention that there is not much difference between choosing particular gauge from the beginning and working with specific gauge-invariant combination.
@@
Instead, the second approach in tackling the gauge issue is by defining a gauge suitable to best describe the observable [31, 33, 32]. In [31], they argue that the suitable gauge is one in which the coordinates are fixed according to the positions of the interferometer mirrors. This is known as the synchronous gauge, or TT gauge. [31] show that the observable GW spectrum of GWs in the TT gauge and Poisson gauge coincide. It is important to note that scalar-induced tensor perturbations can be interpreted as freely propagating gravitational waves—and thus effectively treated as linear, first-order metric perturbations—only once the source term becomes inactive and the modes are well inside the horizon, in the deep sub-horizon regime. In this context, the gauge dependence associated with the perturbations no longer persists, as the gravitational waves evolve independently. A similar argument was made by [32], in which they show that the scalar-induced gravitational waves behave similar when calculated in gauges which are “well-behaved” on small scales. The authors point out that the Newtonian gauge is suitable both for the calculation and the physical interpretation. Further mentioning other gauges, among them the synchronous gauge, to also be “well-behaved” in small scales in a universe dominated by perfect fluid with constant equations of state.
Despite significant progress in previous studies, certain limitations persist. In particular, many approaches impose gauge conditions at an early stage, which can hide leftover gauge effects. While gauge-invariant formulations offer a rigorous framework, they can still present challenges in relating theoretical constructs to measurable quantities. In this review, we present a general framework for deriving the equation of motion and source term for SIGWs without imposing a specific gauge choice. This approach provides a unified and practical framework for evaluating the source term. We examine its behavior across three commonly used gauges—synchronous, Poisson, and uniform curvature—to illustrate how gauge choices influence the resulting expressions. By conducting both analytical and numerical comparisons of the corresponding kernels, we systematically quantify gauge-related discrepancies and clarify the degree to which SIGWs can be considered gauge-invariant at the level of observables.
@@
In Sections 2 and 3, we introduce the general framework for cosmological perturbations of the metric and then derive the perturbed Einstein field equations. In Section 3.3, the source term is presented without imposing a gauge. Further on the gauge transformation and gauge-invariant variables are presented in Section 4. We work with the approaches mentioned to show the behaviour of the scalar-induced gravitational waves in different gauges. In Sections 5 and 6, we study the source-term of the SIGWs and the subsequent kernel in three different gauges.
@@ 2 Cosmological perturbation of the metric
This section introduces the perturbative framework used to study small fluctuations around the homogeneous and isotropic FRW universe. The decomposition into scalar, vector, and tensor modes allows us to trace how different physical effects—such as density fluctuations and gravitational waves—emerge from different types of perturbations. Understanding this structure is fundamental before identifying which modes source second-order gravitational waves.
Any tensorial quantity $A$ can be decomposed into background and inhomogeneous parts. This decomposition is performed in a Friedmann–Lemaître–Robertson–Walker (FLRW) cosmological spacetime, employing a foliation defined by the conformal time $\tau$, such that the hypersurfaces of constant $\tau$ are spatially homogeneous and isotropic. This standard background provides the basis for the perturbative expansion used throughout the paper. The decomposition follows
@@
where $\delta g_{\mu\nu}$ and $\delta^{(2)}g_{\mu\nu}$ is the perturbed metric in first and second order. Using a method known as (3+1) decomposition, also known as scalar–vector–tensor (SVT) decomposition, 3-dimensional quantities are split into scalar, vector and tensor components corresponding to the transformational behavior on spatial hypersurfaces. The scalar, vector and tensor modes are defined as follows [13, 22]:
  • Scalar perturbations: $\phi^{(r)}$, $\psi^{(r)}$, $w^{(r)}$, and $h^{(r)}$,
  • Vector perturbations: transverse $w^{(r)}_{i}$, $h^{(r)}_{i}$,
  • Tensor perturbations: transverse-traceless $h^{(r)}_{ij}$.
Throughout this work, we denote the perturbation order with a superscript in parentheses, e.g., $\phi^{(1)}$, $h^{(2)}_{ij}$. The components of the perturbed spatially flat Robertson - Walker metric can be written as
@@
The vector perturbations are transverse, i.e. $\partial^{i}w_{i}^{(r)}=\partial^{i}h_{i}^{(r)}=0$, and $h_{ij}^{(r)}$ is a symmetric transverse and trace-free tensor, i.e. $\partial^{i}h_{ij}^{(r)}=h_{i}^{i(r)}=0$. Finally, $D_{ij}=\partial_{i}\partial_{j}-\frac{1}{3}\nabla^{2}\delta_{ij}$ is the trace-free operator.
@@
$g^{00}$ $=-a^{-2}(\tau)\left(1-2\phi^{(1)}-\phi^{(2)}+4\left(\phi^{(1)}\right)^{2}-\partial^{i}w^{(1)}\partial_{i}w^{(1)}\right),$
$g^{0i}$ $=a^{-2}(\tau)\left[\partial^{i}w^{(1)}+\frac{1}{2}\left(\partial^{i}w^{(2)}+w^{i(2)}\right)+2\left(\psi^{(1)}-\phi^{(1)}\right)\partial^{i}w^{(1)}\right.$
@@
$g^{ij}$ $=a^{-2}(\tau)\left[\left(1+2\psi^{(1)}+\psi^{(2)}+4\left(\psi^{(1)}\right)^{2}\right)\delta^{ij}+D^{ij}\left(h^{(1)}+\frac{1}{2}h^{(2)}\right)\right.$
@@ 3 Einstein field equations
To determine the dynamics of the perturbations introduced earlier, we derive the Einstein field equations order by order. The key focus is on how scalar perturbations evolve and interact at second order to source tensor modes—i.e., gravitational waves. This derivation lays the groundwork for computing the source term of SIGWs. The Einstein field equation is defined as
$G^{\mu}{{}_{\nu}}=R^{\mu}{{}_{\nu}}-\frac{1}{2}g^{\mu}{{}_{\nu}}R=\kappa^{2}T^{\mu}{{}_{\nu}}.$ (9)
@@ 3.1 Background Einsteins equations
where $\bar{P}$ and $\bar{\rho}$ are the background quantities for pressure and energy-density and $\mathcal{H}=\frac{a^{\prime}}{a}$ is the conformal Hubble parameter [^2] The prime notation indicates derivatives with respect to conformal time $\tau$..
@@ 3.2 First order Einsteins equations
$a^{-2}\left[6\mathcal{H}^{2}\phi^{(1)}+6\mathcal{H}\psi^{(1)^{\prime}}+2\mathcal{H}\nabla^{2}w^{(1)}-2\nabla^{2}\psi^{(1)}-\frac{1}{2}\partial_{k}\partial^{i}D^{k}{{}_{i}}h^{(1)}\right]$ $=-\kappa^{2}\rho^{(1)},$ (11)
@@
$-a^{-2}$ $\left(2\mathcal{H}\phi^{(1)^{\prime}}+\left[4\mathcal{H}^{\prime}+2\mathcal{H}^{2}\right]\phi^{(1)}+4\mathcal{H}\psi^{(1)^{\prime}}+2\psi^{(1)^{\prime\prime}}\right.$
$\left.+\frac{2}{3}\nabla^{2}\left(\phi^{(1)}-\psi^{(1)}+2\mathcal{H}w^{(1)}+w^{(1)^{\prime}}\right)\right)=\kappa^{2}P^{(1)},$ (13)
@@
Here $\rho^{(1)}$, $v^{(1)}_{i}$, and $\pi^{(1)}_{ij}$ are the first-order scalar, vector, and tensor components of the energy-momentum tensor, respectively.
@@ 3.3 Second-order Einstein field equation and source term
For understanding the source-term of scalar-induced gravitational waves, we will be studying the spatial component of second-order Einstein tensor, specifically the transverse, trace-free part. We do this by first applying the projection tensor $P^{li}_{jm}$ to the field equation [34], i.e. $P^{li}_{jm}G^{(2)m}{{}_{l}}=\kappa^{2}P^{li}_{jm}T^{(2)m}{{}_{l}}$. Accordingly, the second-order, scalar, vector and tensor modes are removed. The projected spatial Einstein tensor, Section C.3, and energy-momentum tensor, Section D.3, is presented in Appendix E. Using the relation for the equation of state, $\omega=\bar{P}/\bar{\rho}$, and the background Section 3.1, we can find the wave-equation for the second-order tensor mode,
@@
where the source term is given in Appendix F. The second-order tensor perturbation $h^{(2)}_{ij}$ is sourced by quadratic combinations of first-order scalar modes such as $\phi^{(1)}$, $\psi^{(1)}$, and their spatial derivatives. We decided to find this generic form of the source term Appendix F as this can be simplified to further understand the behavior in different gauges by imposing the gauge condition.
In the source term for SIGWs, only quadratic combinations of first-order scalar perturbations appear, while purely second-order scalar perturbations are absent. This is not an ad hoc assumption but a direct consequence of the structure of the Einstein equations at second order. The key reason lies in how the second-order tensor perturbations are sourced: they arise from the non-linear coupling of first-order scalar modes through products of perturbations that project onto the transverse-traceless (TT) part of the spatial Einstein equations.
Crucially, scalar perturbations—whether first or second order—affect only the scalar sector in the scalar–vector–tensor (SVT) decomposition. When projecting the Einstein equations onto the transverse-traceless (TT) part, only specific quadratic combinations of first-order scalars contribute, as purely second-order scalars do not source tensor modes due to their spin-0 nature [^3] The SVT decomposition classifies perturbations by their behavior under spatial rotations: scalars (spin-0), vectors (spin-1), and transverse-traceless tensors (spin-2). See [34, 13].. This explains why the source term in Eq. 15 includes only first-order combinations.
This structure is a robust feature of second-order cosmological perturbation theory and has been systematically demonstrated in several works, including [14, 35]. Therefore, the form of Eq. 15 is not the result of an additional assumption, but follows from the TT projection of Einstein’s equations and the properties of metric perturbations under SVT decomposition.
@@
where $S_{\lambda}(\tilde{\tau},\mathbf{k})$ includes the Fourier transformation of the source Appendix F and depends on the evolution of the scalar modes. Any scalar mode can be split into a transfer function $T(k\tau)$, which includes the time evolution of the modes after they re-enter the horizon, and the initial condition $\zeta_{\mathbf{k}}$, determined by inflation. For example, the scalar potential $\phi^{(1)}(\mathbf{k},\tau)$ can be split as
$\phi^{(1)}(\mathbf{k},\tau)=\frac{3(1+\omega)}{5+3\omega}T_{\phi}(k\tau)\zeta_{\mathbf{k}}\overset{\omega\rightarrow 1/3}{=}\frac{2}{3}T_{\phi}(k\tau)\zeta_{\mathbf{k}}.$ (18)
@@
where the Green’s function $G_{\mathbf{k}}$ in radiation era is [36]
$kG_{\mathbf{k}}(x,\tilde{x})=\sin(x-\tilde{x})\ ,$ (24)
where $x=k\tau$. Here, we adopt the Green’s function that solves the tensor wave equation in a radiation-dominated background with causal boundary conditions, such that it vanishes for $\tau<\tilde{\tau}$ [36, 35]. The factor of $k$ in the sine function argument arises from the structure of the Fourier-space solution in conformal time. The fraction of GW energy density per logarithmic wavelength is defined by
@@
$\rho_{GW}(\tau)=\frac{1}{16a^{2}(\tau)\kappa^{2}}\overline{\langle(\nabla h_{ij})^{2}\rangle}.$ (26)
The overline above the power spectrum is the oscillation average, derived from the average of the kernel. While the second-order tensor perturbations can be constructed to be gauge-invariant under certain conditions, the derived observable—the spectral energy density of gravitational waves, $\Omega_{GW}$—can still exhibit gauge dependence. This is particularly evident in gauges that are not well-behaved on small scales, such as the comoving slicing gauge, where the induced GW spectrum can differ from that in the Newton gauge. Therefore, care must be taken in choosing an appropriate gauge when interpreting the physical implications of SIGWs [32, 33, 37, 38].